# ACPO: AI-Enabled Compiler-Driven Program Optimization

### v2

@@ 5.2. Combined Results
Sz%: Size bloat, Sp: Speedup. Multicolumns 1, 3, and 4 are wrt. O3 and 2, is ACPO-FI compared against MLGO (Trofin et al., 2021).
| Benchmark | ACPO-FI | wrt. MLGO | ACPO-LU | Combined | ||||
| Sz% | Sp | Sz% | Sp | Sz% | Sp | Sz% | Sp | |
| auto_bitcount | 1 | 0.999 | 1 | 1.005 | 1.001 | 1.003 | 1 | 1.001 |
| auto_qsort1 | 1 | 0.994 | 1 | 1.003 | 1 | 0.982 | 1.91 | 0.989 |
| auto_susan_c | 1 | 0.999 | 1.567 | 1.122 | 3.685 | 1.035 | 19.797 | 1.066 |
| auto_susan_e | 1 | 1.023 | 1.567 | 1.04 | 3.685 | 1.019 | 19.797 | 1.026 |
| auto_susan_s | 1 | 0.98 | 1.567 | 1.008 | 3.685 | 1.038 | 19.797 | 0.995 |
| bzip2d | 1.441 | 0.991 | 1.239 | 1.021 | 2.224 | 1.01 | 3.541 | 1.014 |
| bzip2e | 1.441 | 1.007 | 1.189 | 1.041 | 2.35 | 1.001 | 3.546 | 1.015 |
| cons_jpeg_c | 1.288 | 1.017 | 1.2 | 1.005 | 5.385 | 0.986 | 5.086 | 1.027 |
| cons_jpeg_d | 1.289 | 0.923 | 1.19 | 1.335 | 5.4 | 1.073 | 4.807 | 1.026 |
| cons_lame | 1.437 | 1.027 | 1.137 | 1.007 | 4.501 | 1.03 | 4.729 | 1.034 |
| cons_mad | 1.217 | 1.071 | 1.16 | 1 | 4.321 | 0.99 | 3.193 | 0.968 |
| cons_tiff2bw | 1.18 | 1.002 | 1.351 | 1.23 | 4.502 | 0.943 | 2.471 | 0.987 |
| cons_tiffmedian | 1.18 | 1.017 | 1.048 | 1.056 | 1 | 0.97 | 3.021 | 0.983 |
| network_dijkstra | 1 | 0.993 | 1 | 0.996 | 1 | 0.99 | 1 | 1.037 |
| network_patricia | 1 | 0.98 | 1.052 | 0.99 | 1 | 0.974 | 1 | 1.004 |
| office_ghostscript | 1.275 | 1 | 1.16 | 1.04 | 1 | 1 | 1.12 | 1.001 |
| office_ispell | 0.996 | 1 | 1.17 | 1.02 | 1 | 1 | 1 | 1 |
| office_rsynth | 1 | 1 | 1.12 | 1.04 | 2.024 | 1.01 | 2.024 | 1.011 |
| office_strings | 1 | 1.006 | 1 | 1 | 1 | 1 | 1 | 1 |
| secur_blowfish_d | 1 | 1.002 | 1 | 1.002 | 1 | 0.978 | 1 | 0.965 |
| secur_blowfish_e | 1 | 1.003 | 1.39 | 1.029 | 1 | 1.015 | 1 | 0.981 |
| security_pgp_d | 1.861 | 0.993 | 1.12 | 1 | 1 | 0.991 | 13.144 | 1 |
| security_pgp_e | 1.861 | 1.017 | 1.12 | 1.042 | 1 | 1.006 | 13.144 | 1.006 |
| secur_rijndael_d | 1 | 0.99 | 1.39 | 1.039 | 3.823 | 0.966 | 1 | 0.966 |
| secur_rijndael_e | 1 | 1.001 | 1.09 | 1.035 | 3.823 | 0.974 | 1 | 0.995 |
| security_sha | 0.999 | 1.27 | 1.009 | 1.132 | 1 | 1.174 | 0.999 | 1.501 |
| tele_adpcm_c | 1 | 1 | 0.996 | 1.001 | 1 | 1.008 | 1 | 1.038 |
| tele_adpcm_d | 1 | 0.997 | 1 | 1 | 1 | 1.132 | 1 | 1.138 |
| telecom_CRC32 | 1 | 1.094 | 1 | 1.066 | 1 | 1.042 | 1 | 1.112 |
| telecom_gsm | 1.81 | 1.076 | 1.05 | 1.04 | 1.815 | 1.15 | 1.81 | 1.117 |
| Geomean | 1.152 | 1.021 | 1.161 | 1.041 | 1.818 | 1.01 | 2.442 | 1.024 |
The combined results are shown in Tables 2 and 3. We observe a slight increase in performance benefit for both benchmarks, i.e., from 4% to 4.5% for Polybench and from 2% to 2.4% for Cbench, respectively. Note that we don’t deploy ACPO-FI on Polybench as it consists of small nested-loop kernels which might not directly benefit from an optimized inliner, however, the experimental result reveals that indirectly it can. The underlying reason is that by it can provide more opportunities for the subsequent passes registered under O3 (Ashouri et al., 2022). LLVM’s O3 pipeline has a complex behavior when it comes to pass activation and as mentioned earlier, several passes including Loop Unrolling are activated multiple times thus, it is a hard task to quantify the benefits of each model when deployed together and the accurate measurement of partial benefits, especially when optimizations are working at different scopes, i.e., Call Graph vs Function level.
@@
One exceptionally good performing benchmark is security_sha, which we had already optimized by 27% (show on Table 3) using ACPO’s Inliner, now with 50% performance gain. There are a series of decisions made by ACPO models that we highlight an excerpt under Listing 5. Additionally, each corresponding line at Listing 6 shows the decisions made by LLVM’s O3. Note that, partial and runtime unrolling are both disabled by default at LLVM O3 and thus, F[sha_transform] Loop %for.body5 is not unrolled in Listing 6. To be abundantly fair, we measured the performance of O3 when these two flags were forced enabled as well and we noticed that the decision to unroll the loop was changed to 8 as opposed to ACPOLUModel’s prediction of 16 shown in Listing 5. However, evaluations showed that still ACPO was outperforming O3 by a margin of 7% on unrolling this particular benchmark. Further inspection revealed that O3’s threshold mechanism built into the current version of Loop Unroll pass was limiting the assignment of higher values of unroll count even though the trip count was available at compile time. Similarly, the decision to not inline several callsites including sha_transform is limited by the Inliner’s threshold mechanism.
6. Discussion
@@
ACPO provides a class of features at varying LLVM scopes, therefore, new passes can utilize the already available features with minimal or no efforts. It also provides the flexibility of supporting other types of features, i.e., Graph-based, embeddings, etc. Additionally, we provide the Ahead-Of-Time (AOT) compilation facilities as an alternative method of inference to the developers. As mentioned in Section 3.5, ACPO uses Python named pipes to perform interprocess communication. Although this allows for greater flexibility, it adds a slight overhead to the compilation time (Shown in Table 1). Developers can switch between these modes when building the compiler. The flexibility of choosing between the two modes allows for a fast iterative prototyping of ML models during design and training without the need to rebuild the compiler infrastructure after every change and also, an efficient compilation overhead at deployment once the ML models are optimized and ready to be shipped.
7. Conclusion
@@
In this paper, we presented ACPO, the first scalable and modular ML-guided compiler framework capable of optimizing performance for LLVM. We demonstrated that using ACPO framework, we are able to enable Loop Unroll and Function Inline passes to benefit from a seamless integration of ML models, i.e., ACPOLUModel and ACPOFIModel, for which we experimentally achieved on average 4% and %4.5 on Polybench, and 2.1% and 2.5% on Cbench. ACPO provides a generalized feature collection class for which compiler developers can add to their instantiated ACPOModel class of their need and provides utility classes to use other types of features as well. Future works will be focused on extending ACPO framework by enabling other features and optimization passes inside LLVM’s pipeline and providing innovative software characterization methods using large Languages Models (LLMs).

### v3

@@ V-C Combined Results
An inspection of O3’s pipeline reveals that function inlining triggers before the first instance of loop unrolling, and thus, an enhanced function inlining may benefit the subsequent passes, including loop unrolling. Note that there are a number of passes in between the two instances of our models, i.e., DCE, etc, which change the code size of an IR, and thus, the combined code size bloat does not reflect the addition of the two instances of deployments. The combined results are shown in Tables VI and V. We observe a slight increase in performance benefit for both benchmarks, i.e., from 4% to 4.5% for Polybench and from 2% to 2.4% for Cbench, respectively. Note that we don’t deploy ACPO-FI on Polybench as the benchmark consists of small kernels of nested loops, which might not directly benefit from an optimized inliner. However, the combined experimental result reveals that it can be done indirectly. LLVM’s O3 pipeline has a complex behaviour when it comes to pass activation, and as mentioned earlier, several passes, including Loop Unrolling, are activated multiple times; thus, it is a hard task to quantify the benefits of each model when deployed together and the accurate measurement of partial benefits, especially when optimizations are working at different scopes, i.e., Call Graph vs Function level. One exceptionally good performing benchmark is security_sha, which we had already optimized by 27% (shown on Table V) using ACPOFIModel, now with 50% performance gain. There are a series of decisions made by ACPO models that we highlight an excerpt under Listing 4. Additionally, each corresponding line at Listing 5 shows the decisions made by LLVM’s O3. Note that partial and runtime unrolling are both disabled by default at LLVM O3 and thus, F[sha_transform] Loop %for.body5 is not unrolled in Listing 5. To be abundantly fair, we measured the performance of O3 when these two flags were forced enabled as well and we noticed that the decision to unroll the loop was changed to 8 as opposed to ACPOLUModel’s prediction of 16 shown in Listing 4. However, evaluations showed that still ACPO was outperforming O3 by a margin of 7% on unrolling this particular benchmark. Further inspection revealed that O3’s threshold mechanism built into the current version of Loop Unroll pass was limiting the assignment of higher values of unroll count even though the trip count was available at compile time. Similarly, the decision to not inline several callsites including sha_transform is limited by the LLVM Inliner’s threshold mechanism.
VI Discussion
ACPO is the first scalable framework for bringing ML into compilers. It provides the ability to reproduce models on different architecture very fast and the methodology lays out the steps to regenerate data, retrain a new model, and easily deploy with LLVM. Compiler autotuning and optimization space have benefited from ML for almost two decades. As we pointed out earlier, there are many recent works leveraging ML, however, the goal of ACPO is to provide a novel flexible framework for the LLVM community to leverage ML on any optimization passes of interest with minimal efforts. Nevertheless, there are a number of challenges and shortcomings we would like to mention here.
Integration and dependencies
@@ VI Discussion
ACPO attempts to alleviate the challenging task of supply chain management and the use of licensed tools by separating the ML framework from the compiler. We showed both the integration of PyTorch and TensorFlow among other moving objects, i.e., Autotuner, into our framework, however, upstreaming would remain challenging and requires community efforts to shape the next generations of ML-enabled compilers. ACPO provides a class of features at varying LLVM scopes, therefore, new passes can utilize the already available features with minimal or no effort. It also provides the flexibility of supporting other types of features, i.e., Graph-based, embeddings, etc. Additionally, we provide the Ahead-Of-Time (AOT) compilation facilities as an alternative method of inference to the developers. As mentioned in Section III-D, ACPO uses Python named pipes to perform interprocess communication. Although this allows for greater flexibility, it adds a slight overhead to the compilation time (Shown in Table II). Developers can switch between these modes when building the compiler. The flexibility of choosing between the two modes allows for a fast iterative prototyping of ML models during design and training without the need to rebuild the compiler infrastructure after every change and also, an efficient compilation overhead at deployment once the ML models are optimized and ready to be shipped.
Multi-objective optimization
ACPO optimizes the performance of a code region with intelligent unrolling decisions. However, in a number of cases, we experimentally observe that the emitted code size is also increased. This is a(n) (in)direct result of Inliner and Loop Unroll which increase the number of instructions and thus, the code size. Although this is outside the scope of the current work, an enhanced version may also benefit from taking into account both of the objectives in its exploration strategy by identifying the Pareto optimality [58, 41, 59, 60]. Similar strategies are employed for a number of adjacent problems; i.e, Multicore Embedded Systems [61, 62] and Compiler autotuning [58, 41, 59].
VII Conclusion
@@
In this paper, we presented ACPO, a scalable and modular ML-guided compiler framework capable of optimizing performance for LLVM. We demonstrated that using the ACPO framework, we are able to enable Loop Unroll and Function Inline passes to benefit from a seamless integration of ML models, i.e., ACPOLUModel and ACPOFIModel, for which we experimentally achieved on average 4% and %4.5 on Polybench, and 2.1% and 2.5% on Cbench. ACPO provides a generalized feature collection class for which compiler developers can add to their instantiated ACPOModel class of their need and provides utility classes to use other types of features as well.